# Evaluating Meta-Regression Techniques: A Simulation Study on Heterogeneity in Location and Time

### v1

@@ 5.3.1 Impact of Small versus Large Time Effect
Between all methodology specifications, the standard methodology performed best when faced with a large time effect, but when faced with a small time effect $FE_{l,Trend}$ and $FE_{lt}$ performed just as well as the standard methodology. Figures 7 - 9 in Appendix C show performance criteria from a sample result that illustrates this point (see the supplementary materials for the full results). Figure 7 in Appendix C shows that when time heterogeneity is larger, power is substantially worse for the location-level methodology specifications $ME_{l}$, $RE_{l}$ $FE_{l}$, power is equally poor for $FE_{t}$ and power remains high and unchanged for the remaining specifications. The residuals MSE is accordingly worse for the location-level specifications as seen in Figure 8 in Appendix C. Figure 9 also in Appendix C shows that in the face of higher time heterogeneity, the accuracy of the trend estimate slightly worsens, but its precision stays the same. These results support the first hypothesis since $ME_{l}$, $RE_{l}$, $FE_{l}$ $FE_{t}$ did not perform as well as the other specifications that jointly control for time and location variation. This also means that the results do not support the second hypothesis. Note that when time effects are large, the $FE_{l,Trend}$ $FE_{lt}$ specifications performed second best. Time heterogeneity was controlled for with $FE_{t}$ $FE_{lt}$, and measured with $FE_{s,Trend}$ $FE_{l,Trend}$. The first two control for time with fixed effects, while the last two both control and capture the effect of any time trends shared by all locations in the meta-regression. Immediately, the results of $FE_{t}$ show that time-level fixed effects should not be used for data similar to the simulated data since this methodology has substantially less power compared to all the other methodologies tested for the twelve cases. Consequently, the residuals MSE also performed substantially worse. These problems were even worse in the cases where large time effects were present, so $FE_{t}$ $FE_{lt}$ perform even worse than when the time effects are smaller. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
@@
The trend for the two methodology specifications with a trend ($FE_{s,Trend}$ $FE_{l,Trend}$) have a rescaled trend term of $[-1,-0.5,0,0.5,1]$ when $t=5$ $n=5$ (where $t$ is the number of time periods and $n$ is the number of locations). When $n=9$ $t=5$, the rescaled trend term is $[-1,-0.75,-0.5,-0.25,0]$, while $n=15$ $t=5$ the rescaled trend term is $[-1,-0.86,-0.71,-0.57,-0.43]$. The trends’ sample size is so small because this simulation includes 5, 9, or 15 locations, and each location has 5 studies over 5 years. The two specifications with this trend ($FE_{s,Trend}$ $FE_{l,Trend}$) both perform very well in terms of power and the residuals MSE. However, the key difference is that $FE_{s,Trend}$ provides very inaccurate estimates for the trend coefficient, while $FE_{l,Trend}$ has almost perfect accuracy for the trend coefficient. Figure 9 in Appendix C shows a sample result that illustrates exactly this point. Only when there is a large time effect is the power slightly worse for $FE_{l,Trend}$ (compared to a small time effect). This causes the residuals MSE and the estimate of the trend coefficient to be slightly worse (but still far better than $FE_{s,Trend}$). Note that all specifications perform slightly worse when the time effect is large, supporting the third hypothesis. When large time effects are present the performance also suffers in the same way for $RE_{l}$ $FE_{l}$. Note that the estimator bias is very small for all cases and methodology specifications (but note that the estimator for $FE_{t}$ has a much larger variance).
5.3.2 Impact of Small versus Large Location Effect
@@
When the location effect is large and the number of locations is held constant, all methodology specifications performed just as well (excluding $FE_{t}$ which faced worse power, and $FE_{s,Trend}$ where the trend estimate worsened). Note that these specifications performed poorly in Sections 5.3 and 5.3.1 though, so these differences are trivial. Figures 10 to 11 in Appendix C illustrate this result since power, estimator bias, and the residuals MSE are unchanged. Note that Figure 12 in the same Appendix C shows that in the face of higher location heterogeneity, both the precision and accuracy of the trend estimate stay the same. If there are no confounding factors, this means that no matter how large the differences are between each location in the studies within meta-regression data, either joint and non-joint location-level heterogeneity methodology specifications do well in controlling location-level heterogeneity. This would also suggest that location-level heterogeneity can be better controlled compared to time-level heterogeneity. The third hypothesis contradicts with this result though based on the common finding in the meta-analysis simulation literature and accordingly calls for further analysis to ensure there are no confounding factors. To support the above hypothesis, further testing that ensures the given scale of location heterogeneity matches the given scale of time heterogeneity would need to be done.
@@ 5.3.3 Impact of Number of Locations
In the face of more locations, the worst performing methodology specifications were all the non-joint specifications ($FE_{t}$, $FE_{l}$, $RE_{l}$ $ME_{l}$). This is observed in Figure 14 in Appendix C since the residuals MSE is a little bit worse for those specifications. Note that Figure 13 in Appendix C shows that when there are more locations power for all methodology specifications raises a little bit higher. Lastly, Figure 15 in the same in Appendix C shows that when there are more locations, the accuracy of the trend estimate slightly worsens but its precision decreases (supporting the third hypothesis). Sample size is likely the key to performance improvements since increasing from 5 to 9 countries almost doubles the sample size, and increasing from 5 to 15 countries triples the sample size (since we design each location to have an identical sample size). Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion. The only specification that faced accuracy problems was the estimator for $FE_{t}$, which although it had very low bias it also had a much larger variance. If there are no confounding factors, this means that the more diverse the studies are within a meta-analysis, the stronger the estimators will be (as long as the joint heterogeneity is appropriately accounted for in the methodology specification).
@@ 5.3.4 Summary of Results
Since there are many parameters and cases to capture all the nuances of meta-regression data sets, we look at each parameter and case ceteris paribus. First, Section 5.3 shows that all methodology specifications perform better when more covariates are present (only excluding $RE_{l}$). This is a common result in the meta-analysis simulation literature, and was found by the uniform decreased MSE of the residual. Second, Section 5.3.1 shows that in the face of a large time effect the standard methodology performed best, second best were the $FE_{lt}$ and the newly posed $FE_{l,Trend}$ joint methodologies, and lastly the worst performing were the remaining non-joint methodologies. This supports the first hypothesis. Note that the accuracy of trend estimates lower but precision increases. Third, Section 5.3.2 shows that in the face of a large location effect, joint and non-joint specifications perform just as well. This does not support the first or third hypotheses, unlike all other scopes reported in the subsections of Section 5.3 (i.e. covariates, time effects, and number of locations). Lastly, Section 5.3.3 shows two differences when the number of locations change. One difference is that all methodology specifications have more power, and the variance of the estimator is smaller. This means that any model will perform better when more locations are present, likely because the sample size of the respective meta-regression will be larger. Another difference is that for $FE_{l,Trend}$, the variance of the trend estimate improves, but the precision worsens. The precision only worsens marginally in this simulation, but caution should be used regardless when there are additional locations in a meta-regression. These two differences support the third hypothesis. Tables 5 & 6 in Appendix E summarize how increases or decreases in each type of heterogeneity can affect performance criteria as increases, decreases or little to no change. Note that these tables solely indicate how well each methodology specification handles larger heterogeneity compared to smaller heterogeneity, these should not be used as indicators of methodology performance. Overall, in all of these scopes we can see the standard methodology performed best, and any of the joint methodology specifications performed a very close second best (including $FE_{lt}$ and the newly introduced $FE_{l,Trend}$, but excluding $FE_{s,Trend}$ due to poor trend estimates). This also supports the first hypothesis, but does not support the second hypothesis. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
6 Example Applications & Parameter Extraction for Simulations
@@
As mentioned in Section 5.1, these parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from said data. Most studies in the meta-analysis simulation literature extract parameters from one particular meta-regression to ensure the consistency of parameters (Pereira et al., 2010; Schwarzer et al., 2024; Dormann et al., 2020). Before discussing parameter extraction, this section will review where these meta-regressions can be applied. Meta-regressions are most commonly used to analyze the consistency of medical laboratory study results, but they are also regularly used in many other fields of study as well. Section 2 showed that meta-regressions are used in experimental economics data sets to evaluate the consistency of behavioral laboratory study results. Note that data in this field typically observe variations in both the number of locations and the number of static time periods per study. Meta-regressions have also been used to analyze the consistency of various types of financial data including market data, banking data, and corporate data (Buch et al., 2018; Guo et al., 2023; Dao & Ta, 2020). Note that this body literature uses data that observe variations in both the number of locations and the number of time periods. If there are studies that have multiple time periods within, these need to be handled with a different methodology specification that additionally accounts for time series within each study. The most notable meta-analysis in the finance literature is a collaboration between the central banks of 17 countries, where each country has time-series data (Buch et al., 2018). Subfields of economics that measure willingness to pay and contingent valuation (such as natural resource and environmental economics) also use meta-regressions to assess the consistency of these measures across similar studies (Wehkamp et al., 2018; Vesco et al., 2020; Afesorgbor and Demena, 2022; Nelson & Kennedy, 2009).
@@
There are steps any practitioner can follow to extract parameters from their data, in order to conduct simulations to assess which methodology specification would perform best for their data. With the extracted means, variances, and covariances, we can then generate synthetic data following the same simulation design described in Section 5.1.2. Note that this assumes that the data is normally distributed and independent, if either are not true estimators will perform very poorly as discussed in Section 5.1 (Blázquez-Rincón et al., 2023; McGrath et al., 2023; Bohnet, 2007; Bergtold et al., 2010; Poudyal & Spanos, 2022; Rubio‐Aparicio et al., 2018). The data can then be modeled using a Monte Carlo simulation using the same performance criteria defined in Section 5.2, but now we also have the option to compare the simulation results with the findings of the meta-analysis. This provides value in two ways: first, this can also be used to assess the replicability of results using linked simulation data. Second, the performance of the initial methodology specification can be compared with others to dictate which specification is optimal for the data in question (as illustrated in the previous sections simulation). Note that before doing either of these, a baseline needs to be established by ensuring consistency between model results from the real data and the linked simulation data. This extension can be applied with ease to simulation conducted in Section 5, and both of the location and time effects could also easily be calibrated.
@@
    • Any missing data needs to be handled in whichever way is most appropriate (Little & Rubin, 2019; Baraldi and Enders, 2010; Graham, 2009; Buuren, 2018; J. Lee & Beretvas, 2023).
@@ 7 Conclusion
The purpose of this simulation study is to determine which meta-regression methodology performs optimally when joint location and time heterogeneity are present between studies within a meta-regression. This provides three contributions that are essential for both the meta-regression simulation literature, and practitioners in any discipline that use meta-regressions. First, the standard methodology only controls for heterogeneity at the study level, so we test the performance of seven methodology specifications new to meta-regressions in the face of varying degrees of location and time level heterogeneity. Second, the efficacy of the newly proposed location-level fixed effects in the experimental economics literature and study-level fixed effects with a trend specifications in the meta-regression simulation literature are tested. Third, steps have been outlined for any practitioner to select the optimal model for their data. This is essential to ensure that the selected model matches the type of heterogeneity in the data. Our simulation embeds both location and time heterogeneity in the dependent variable and tests various parameter combinations (sample size, number of covariates, number of locations and magnitude of location and time effects). We evaluate ten methodology specifications, including the standard study-level random, fixed, and mixed effects, using performance criteria such as statistical power, estimator bias and confidence interval precision.
@@
Results show that the standard methodology performed best, closely followed by methodologies which jointly control the location and time heterogeneity. Specifications that controlled for only one type of heterogeneity performed worse, including the location-level fixed effects proposed in experimental economics. Study-level fixed effects with a trend produced inaccurate trend estimates, while time-level fixed effects suffered from poor power. Across all methodology specifications, performance improved with more covariates and more locations, likely due to increased sample size. Importantly, when time effects were small, most specifications performed equally well, but when time or location effects were large, performance declined (especially for location-level fixed and random effects). These findings emphasize the importance of selecting models depending on the characteristics of the data, instead of selecting study-level random effects by default. The next subsection provides recommendations practitioners can use as a guide for model selection.
7.1 Recommended Modeling for Practitioners
Before selecting a model, there are three essential things a practitioner needs to do. First, review the meta-regression data to ensure it is ready to model. For instance, there should be no data overlap across studies. If there is, disentangle it, but if there is no possible way to disentangle it, then consider the approach in the literature (Bom & Rachinger, 2020). In addition, if there are any extreme values from rare events, these need to be handled with either controlling for these outliers or other processes described in the literature (Günhan et al., 2020; Thom et al., 2020). Also, if there is any missing data, then that needs to be handled in whichever way is most appropriate for your data. The most common ways to handle missing data are to either drop any rows with missing values, exclude variables that have too many missing values, or impute missing values using a variety of methods (Little & Rubin, 2019; Baraldi and Enders, 2010; Graham, 2009; Buuren, 2018; J. Lee & Beretvas, 2023). Note that each approach to handling missing data has unique tradeoffs  [^3] Dropping any rows with missing values ensures potential explanatory power is not lost in the model, but there is also potentially large losses in data. On the other hand, excluding variables with missing values assures data loss will not be as large, but potential explanatory power will be lost in the model. As for imputing missing values this ensures data and explanatory power will not be lost, but there is the risk that the imputed data is not accurate.. Second, test for heterogeneity in location as well as time, even if one does not suspect the presence of this in the data. Before conducting statistical tests, you can first perform a quick visual test on your data as follows. Location heterogeneity can be checked by calculating the summary statistics of the data by each location to assess if there are differences (or not). Time heterogeneity can be checked by comparing graphs over time per location to determine if any noticeable trends exist. To formally support this, there are statistical tests that can also be conducted to detect heterogeneity including a $\chi^{2}$ test, Breusch–Pagan test, or $I^{2}$ test specific to meta-analyses (Wooldridge, 2010; Cameron & Trivedi, 2005; Pereira et al., 2010; Borenstein et al., 2009). These tests need be used with care though due to their key limitations such as the vague interpretation of $I^{2}$ values, and the Breusch-Pagan test only captures linear heterogeneity for normally distributed errors. Also note that this manuscript and its recommendations only cover joint location and time heterogeneity. So, if your data does not match this, you will likely need to use a different model to better match the heterogeneity present in your data. Third, confirm your data is $NIID$. The meta-analysis simulation literature has shown that when data is not $NIID$ there are issues with the accuracy of estimation, which invalidate any inferences that are made (McGrath et al., 2019, 2023; Blázquez-Rincón et al., 2023; Pereira et al., 2010; Spanos & McGuirk, 2001). Once both are set, then we would recommend either $RE_{s}$, $FE_{s}$, or the newly posed $FE_{l,Trend}$. The data will dictate which of these three methodology specifications are most appropriate. For instance, does the random effects assumption hold? If not, that immediately excludes $RE_{s}$. Also, do you want to estimate the overall time effect instead of just controlling for it? If so, then only the newly introduced $FE_{l,Trend}$ would be appropriate. The next subsection discusses the detailed plans for future research.
7.2 Future Study
@@
Two key areas of future study are being developed to further this research. First, we plan to test the performance of specifications that control only location when only location heterogeneity is present. In the face of joint heterogeneity, the present manuscript found concerns with the location fixed effects methodology specification posed in the experimental economics literature, so this will test if this methodology is appropriate when there is only heterogeneity in location. Second, we plan to test the ten meta-regression methods we tested with machine learning approaches to both perform a robustness check and indicate potential improvements for each method. In the meta-analysis literature, machine learning has only been used to optimize the collection of studies used in a meta-analysis, thus expanding the use to model selection for analysis will be a novel contribution.
Appendix A
@@
Appendix C - Result Figures

### v2

@@ 5.3.1 Impact of Smaller versus Larger Time Effect
Between all methodology specifications, the standard methodology performed best when faced with a large time effect, but when faced with a small time effect $FE_{l,Trend}$ and $FE_{lt}$ performed just as well as the standard methodology. Figures 8 - 11 in Appendix D show performance criteria from a sample result that illustrates this point. Figure 8 in Appendix D shows that when time heterogeneity is larger, power is substantially worse for the location-level methodology specifications $ME_{l}$, $RE_{l}$ $FE_{l}$, power is equally poor for $FE_{t}$ and power remains high and unchanged for the remaining specifications. The estimator bias, variance and MSE are accordingly worse for the location-level specifications as seen in Figures 10 & 11 in Appendix D. Figure 9 also in Appendix D shows that in the face of higher time heterogeneity, the bias in the biased trend estimate for $FE_{s,Trend}$ remains unchanged while the bias in the unbiased trend estimate for $FE_{l,Trend}$ slightly increases. Precision is unchanged for both specifications. These results support the first hypothesis since $ME_{l}$, $RE_{l}$, $FE_{l}$ $FE_{t}$ did not perform as well as the other specifications that jointly control for time and location variation. This also means that the results do not support the second hypothesis. Note that when time effects are large, the $FE_{l,Trend}$ $FE_{lt}$ specifications performed second best. Time heterogeneity was controlled for with $FE_{t}$ $FE_{lt}$, and measured with $FE_{s,Trend}$ $FE_{l,Trend}$. The first two control for time with fixed effects, while the last two both control and capture the effect of any time trends shared by all locations in the meta-regression. Immediately, the results of $FE_{t}$ show that time-level fixed effects should not be used for data similar to the simulated data since this methodology has substantially less power compared to all the other methodologies tested for the twelve cases. Consequently, the estimator bias, variance, and MSE perform substantially worse. These criteria perform worse when large time effects were present, the performance especially suffered for $FE_{t}$ $FE_{lt}$. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
@@
The trend for the two methodology specifications with a trend ($FE_{s,Trend}$ $FE_{l,Trend}$) have a rescaled trend term of $[-1,-0.5,0,0.5,1]$ when $t=5$ $n=5$ (where $t$ is the number of time periods and $n$ is the number of locations). When $n=9$ $t=5$, the rescaled trend term is $[-1,-0.75,-0.5,-0.25,0]$, while $n=15$ $t=5$ the rescaled trend term is $[-1,-0.86,-0.71,-0.57,-0.43]$. The trends’ sample size is so small because this simulation includes 5, 9, or 15 locations, and each location has 5 studies over 5 years. The two specifications with this trend ($FE_{s,Trend}$ $FE_{l,Trend}$) both perform very well in terms of power. However, the key difference is that $FE_{s,Trend}$ provides very inaccurate estimates for the trend coefficient, while $FE_{l,Trend}$ has almost perfect accuracy for the trend coefficient. Figure 9 in Appendix D shows a sample result that illustrates exactly this point. Only when there is a large time effect is the power slightly worse for $FE_{l,Trend}$ (compared to a small time effect). This causes the estimate of the trend coefficient to be slightly worse (but still far better than $FE_{s,Trend}$). Note that all specifications perform slightly worse when the time effect is large, supporting the third hypothesis. When large time effects are present, the performance also suffers in the same way for $RE_{l}$ $FE_{l}$. Note that the estimator bias is very small for all cases and methodology specifications, but note that the estimator for $FE_{t}$ has a much larger variance.
5.3.2 Impact of Smaller versus Larger Location Effect
@@
When the location effect is large and the number of locations is held constant, all methodology specifications performed just as well (excluding $FE_{t}$ which faced worse power, and $FE_{s,Trend}$ where the trend estimate worsened). Note that these specifications performed poorly in Sections 5.3 and 5.3.1 though, so these differences are trivial. Figures 12, 14 & 15 in Appendix D illustrate this result since power, estimator bias, and estimator variance are unchanged. Note that Figure 13 in the same Appendix D shows that in the face of higher location heterogeneity, the already biased trend estimate for $FE_{s,Trend}$ explodes while the unbiased trend estimate for $FE_{l,Trend}$ remains unbiased and its precision increases. If there are no confounding factors, this means that no matter how large the differences are between each location in the studies within meta-regression data, both joint and non-joint location-level heterogeneity methodology specifications do well in controlling location-level heterogeneity (excluding the specifications $FE_{t}$ $FE_{s,Trend}$). Note that the third hypothesis contradicts with this result based on the common finding in the meta-analysis simulation literature. Further analysis is needed to ensure there are no confounding factors and that the scale of location heterogeneity is not too small.
@@ 5.3.3 Impact of Number of Locations
Unexpectedly, in the face of more locations non-joint methodology specifications performed just as well on average as joint methodology specifications. The only specifications that had issues in performance were $FE_{t}$ due to power and $FE_{s,Trend}$ due to bias in the trend estimator. Power is shown in Figure 16 in Appendix D, when there are more locations power for all methodology specifications raises a bit higher. Figure 17 in the same in Appendix D shows that when there are more locations, the already biased trend estimate for $FE_{s,Trend}$ remains unchanged in both bias and precision while bias and precision slightly increase for the unbiased trend estimate for $FE_{l,Trend}$. Sample size is likely the key to performance improvements since increasing from 5 to 9 countries almost doubles the sample size, and increasing from 5 to 15 countries triples the sample size (since we design each location to have an identical sample size). Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion. The only specification that faced accuracy problems was the estimator for $FE_{t}$, which although it had very low bias it also had a much larger variance. If there are no confounding factors, this means that the more diverse the studies are within a meta-analysis, the stronger the estimators will be (as long as the joint heterogeneity is appropriately accounted for in the methodology specification).
@@ 5.3.4 Summary of Results
Since there are many parameters and cases to capture all the nuances of meta-regression data sets, we look at each parameter and case ceteris paribus. First, Section 5.3 shows that all methodology specifications perform better when more covariates are present (only excluding $RE_{l}$). This is a common result in the meta-analysis simulation literature, and was found by the uniform increase in power of each covariate. Second, Section 5.3.1 shows that in the face of a large time effect the standard methodology performed best, second best were the $FE_{lt}$ and the newly posed $FE_{l,Trend}$ joint methodologies, and lastly the worst performing were the remaining non-joint methodologies. This supports the first hypothesis. Note that the accuracy of trend estimates lower but precision increases. Third, Section 5.3.2 shows that in the face of a large location effect, joint and non-joint specifications perform just as well (excluding $FE_{t}$ and $FE_{s,Trend}$. This does not support the first or third hypotheses, unlike all other scopes reported in the subsections of Section 5.3 (i.e. number of covariates, time effects, and number of locations). Lastly, Section 5.3.3 shows two differences when the number of locations change. One difference is that all methodology specifications have more power, and the variance of the estimator is smaller. This means that any model will perform better when more locations are present, likely because the sample size of the respective meta-regression will be larger. Another difference is that for $FE_{l,Trend}$, the variance of the trend estimate improves, but the precision worsens. The precision only worsens marginally in this simulation, but caution should be used regardless when there are additional locations in a meta-regression. These two differences support the third hypothesis. Tables 5 & 6 in Appendix E summarize how increases or decreases in each type of heterogeneity can affect performance criteria as increases, decreases or little to no change. Note that these tables solely indicate how well each methodology specification handles larger heterogeneity compared to smaller heterogeneity, these should not be used as indicators of methodology performance. Overall, in all of these scopes we can see the standard methodology performed best, and any of the joint methodology specifications performed a very close second best (including $FE_{lt}$ and the newly introduced $FE_{l,Trend}$, but excluding $FE_{s,Trend}$ due to poor trend estimates). This also supports the first hypothesis, but does not support the second hypothesis. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
6 Practitioners Guide for Model Selection
Before selecting a model, there are three essential things a practitioner needs to do. First, review the meta-regression data to ensure it is ready to model. For instance, there should be no data overlap across studies. If there is, disentangle it, but if there is no possible way to disentangle it, then consider the approach in the literature (Bom & Rachinger, 2020). In addition, if there are any extreme values from rare events, these need to be handled with either controlling for these outliers or other processes described in the literature (Günhan et al., 2020; Thom et al., 2020). Also, if there is any missing data, then that needs to be handled in whichever way is most appropriate for your data. The most common ways to handle missing data are to either drop any rows with missing values, exclude variables that have too many missing values, or impute missing values using a variety of methods (J. Lee & Beretvas, 2023; Little & Rubin, 2019; Buuren, 2018; Baraldi and Enders, 2010; Graham, 2009). Note that each approach to handling missing data has unique tradeoffs  [^3] Dropping any rows with missing values ensures potential explanatory power is not lost in the model, but there is also potentially large losses in data. On the other hand, excluding variables with missing values assures data loss will not be as large, but potential explanatory power will be lost in the model. As for imputing missing values this ensures data and explanatory power will not be lost, but there is the risk that the imputed data is not accurate.. Second, test for heterogeneity in location as well as time, even if one does not suspect the presence of this in the data. Before conducting statistical tests, you can first perform a quick visual test on your data as follows. Location heterogeneity can be checked by calculating the summary statistics of the data by each location to assess if there are differences (or not). Time heterogeneity can be checked by comparing graphs over time per location to determine if any noticeable trends exist. To formally support this, there are statistical tests that can also be conducted to detect heterogeneity including a $\chi^{2}$ test, Breusch–Pagan test, or $I^{2}$ test specific to meta-analyses (Wooldridge, 2010; Pereira et al., 2010; Borenstein et al., 2009; Cameron & Trivedi, 2005; Baltagi, 2005). These tests need be used with care though due to their key limitations such as the vague interpretation of $I^{2}$ values, and the Breusch-Pagan test only captures linear heterogeneity for normally distributed errors. Also note that this manuscript and its recommendations only cover joint location and time heterogeneity. So, if your data does not match this, you will likely need to use a different model to better match the heterogeneity present in your data. Third, confirm your data is $NIID$. The meta-analysis simulation literature has shown that when data is not $NIID$ there are issues with the accuracy of estimation, which invalidate any inferences that are made (McGrath et al., 2023; Blázquez-Rincón et al., 2023; McGrath et al., 2019; Pereira et al., 2010; Spanos & McGuirk, 2001). Once both are set, then we would recommend either $RE_{s}$, $FE_{s}$, or the newly posed $FE_{l,Trend}$. The data will dictate which of these three methodology specifications are most appropriate. For instance, does the random effects assumption hold? If not, that immediately excludes $RE_{s}$ $ME_{s}$. Also, do you want to estimate the overall time effect instead of just controlling for it? If so, then only the newly introduced $FE_{l,Trend}$ would be appropriate. The next subsection discusses the detailed plans for future research.
7 Applications & Parameter Extraction for Simulations
@@
As mentioned in Section 5.1, these parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from said data. Most studies in the meta-analysis simulation literature extract parameters from one particular meta-regression to ensure the consistency of parameters (Schwarzer et al., 2024; Dormann et al., 2020; Pereira et al., 2010). Before discussing parameter extraction, this section will review where these meta-regressions can be applied. Meta-regressions are most commonly used to analyze the consistency of medical laboratory study results, but they are also regularly used in many other fields of study as well. Section 2 showed that meta-regressions are used in experimental economics data sets to evaluate the consistency of behavioral laboratory study results. Note that data in this field typically observe variations in both the number of locations and the number of static time periods per study. Meta-regressions have also been used to analyze the consistency of various types of financial data including market data, banking data, and corporate data (Guo et al., 2023; Dao & Ta, 2020; Buch et al., 2018). Note that this body literature uses data that observe variations in both the number of locations and the number of time periods. If there are studies that have multiple time periods within, these need to be handled with a different methodology specification that additionally accounts for time series within each study. The most notable meta-analysis in the finance literature is a collaboration between the central banks of 17 countries, where each country has time-series data (Buch et al., 2018). Subfields of economics that measure willingness to pay and contingent valuation (such as natural resource and environmental economics) also use meta-regressions to assess the consistency of these measures across similar studies (Afesorgbor and Demena, 2022; Vesco et al., 2020; Wehkamp et al., 2018; Nelson & Kennedy, 2009).
@@
There are steps any practitioner can follow to extract parameters from their data, in order to conduct simulations to assess which methodology specification would perform best for their data. With the extracted means, variances, and covariances, we can then generate synthetic data following the same simulation design described in Section 5.1.2. Note that this assumes that the data is normally distributed and independent, if either are not true estimators will perform very poorly as discussed in Section 5.1 (Blázquez-Rincón et al., 2023; McGrath et al., 2023; Poudyal & Spanos, 2022; Rubio‐Aparicio et al., 2018; Bergtold et al., 2010; Bohnet, 2007). The data can then be modeled using a Monte Carlo simulation using the same performance criteria defined in Section 5.2, but now we also have the option to compare the simulation results with the findings of the meta-analysis. This provides value in two ways: first, this can also be used to assess the replicability of results using linked simulation data. Second, the performance of the initial methodology specification can be compared with others to dictate which specification is optimal for the data in question (as illustrated in the previous sections simulation). Note that before doing either of these, a baseline needs to be established by ensuring consistency between model results from the real data and the linked simulation data. This extension can be applied with ease to simulation conducted in Section 5, and both of the location and time effects could also easily be calibrated.
@@
    • Any missing data needs to be handled in whichever way is most appropriate (J. Lee & Beretvas, 2023; Little & Rubin, 2019; Buuren, 2018; Baraldi and Enders, 2010; Graham, 2009).
@@ 8 Conclusion
The purpose of this simulation study is to determine which meta-regression methodology performs optimally when joint location and time heterogeneity are present between studies within a meta-regression. This provides three contributions that are essential for both the meta-regression simulation literature, and practitioners in any discipline that use meta-regressions. First, the standard methodology only controls for heterogeneity at the study level, so we test the performance of seven methodology specifications new to meta-regressions in the face of varying degrees of location and time level heterogeneity. Second, the efficacy of the newly proposed location-level fixed effects in the experimental economics literature and study-level fixed effects with a trend specifications in the meta-regression simulation literature are tested. Third, steps have been outlined for any practitioner to select the optimal model for their data. This is essential to ensure that the selected model matches the type of heterogeneity in the data. Our simulation embeds both location and time heterogeneity in the dependent variable and tests various parameter combinations (sample size, number of covariates, number of locations and magnitude of location and time effects). We evaluate ten methodology specifications, including the standard study-level random, fixed, and mixed effects, using performance criteria such as statistical power, estimator bias, estimator variance and confidence interval precision.
@@
Results show that the standard methodology performed best, closely followed by methodologies which jointly control the location and time heterogeneity. Specifications that controlled for only one type of heterogeneity performed worse, including the location-level fixed effects proposed in experimental economics. Study-level fixed effects with a trend produced inaccurate trend estimates, while time-level fixed effects suffered from poor power. Across all methodology specifications, performance improved with more covariates and more locations, likely due to increased sample size. Importantly, when time effects were small, most specifications performed equally well, but when time or location effects were large, performance declined (especially for location-level fixed and random effects). These findings emphasize the importance of selecting models depending on the characteristics of the data, instead of selecting study-level random effects by default.
@@
There are two key areas of future study currently being developed to further this research. First, we plan to test the performance of specifications that control only location when only location heterogeneity is present. In the face of joint heterogeneity, the present manuscript found concerns with the location fixed effects methodology specification posed in the experimental economics literature, so this will test if this methodology is appropriate when there is only heterogeneity in location. Second, we plan to test the ten meta-regression methods we tested with machine learning approaches to both perform a robustness check and indicate potential improvements for each method. In the meta-analysis literature, machine learning has only been used to optimize the collection of studies used in a meta-analysis, thus expanding the use to model selection for analysis will be a novel contribution.
Appendix -Tables
Parameter Symbol Values
Sample size (per study in meta-regression) $n$ 50, 100, 150
Number of covariates $k$ 1, 3, 5
Number of locations (in meta-regression) $n_{l}$ 5, 9, 15
Number of studies (in meta-regression) $n_{s}$ 25, 45, 75
Case Locations Time Effect Location Effect
1 5 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-2,-1,0,1,2\}$
2 5 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-10,-5,0,5,10\}$
3 9 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-2,-1.5,-1,0.5,0,0.5,1,1.5,2\}$
6 15 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10)\}$
7 5 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-2,-1,0,1,2\}$
8 5 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-10,-5,0,5,10\}$
9 9 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-2,-1.5,-1,0.5,0,0.5,1,1.5,2\}$
12 15 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10)\}$
Power Estimator Bias Estimator Variance
$\uparrow n$uparrow $\downarrow$ $\downarrow$
$\downarrow n$downarrow $\uparrow$ $\uparrow$
$\uparrow$ Covariates $\uparrow$ $\downarrow$ $\downarrow$
$\downarrow$ Covariates $\downarrow$ $\uparrow$ $\uparrow$
Appendix A - Survey of Meta-Analyses
@@
Appendix C - Multivariate Data Generation Process
Three Covariate Data Generation Process: This illustrates how each variable is generated and how the true values of the estimators are generated.
Five Covariate Data Generation Process: This illustrates how each variable is generated and how the true values of the estimators are generated.
Appendix D - Result Figures